Juan, Phil, Miguel, and Kim compare their earnings for one month. Phil earned 400 less than twice as much as Juan. Miguel earned one-half as much as Phil. Kim earned three times as much as Juan. Their combined total earnings are $340 less than twice what Kim earned. How much did each person earn and what was the total combined earnings?
J=x
P=2x-400
M=1/2(2x-400)
K=3x
Having trouble on "Their combined total...less than twice what Kim earned."

Answer by lwsshak3(11628)

Juan, Phil, Miguel, and Kim compare their earnings for one month. Phil earned 400 less than twice as much as Juan. Miguel earned one-half as much as Phil. Kim earned three times as much as Juan. Their combined total earnings are $340 less than twice what Kim earned. How much did each person earn and what was the total combined earnings?
***
let x=amt Juan earned
2x-400=amt Phil earned
(2x-400)/2=x-200=amt Miguel earned
3x=amt Kim earned
..
x+2x-400+x-200+3x=2*3x-340
7x-600=6x-340
7x-6x=260
x=260
2x-400=120
x-200=60
3x=780
6x-340=1220
amt Juan earned=$260
amt Phil earned=$120
amt Miguel earned=$60
amt Kim earned=$780
total combined earnings=$1220